# Approximability of Deep Computations

### v3

@@ 6.2.1. Relative compacta of continuous layer transitions
For any topological space $X$, let $\mathrm{C}_{\mathrm{p}}(X)\subseteq\mathbb{R}^{X}$ be the set of all continuous real functions on $X$, endowed with the relative (subspace) topology of the product $\mathbb{R}^{X}$, i.e., the topology of point-wise convergence at each $x\in X$. More generally, given two spaces $X,Y$, the space $\mathrm{C}_{\mathrm{p}}(X;Y)$ is the subspace of the product $Y^{X}=\prod_{x\in X}Y$ consisting of continuous maps $X\to Y$. (“$\mathrm{C}_{\mathrm{p}}$” means “pointwise topology on continuous functions”.)
@@
A topological space $Y$ is angelic if (i) every relatively countably compact subset $A\subseteq Y$ is relatively compact, and (ii) the closure $\overline{A}\subseteq Y$ of any such (relatively compact) $A$ consists precisely of limits of sequences in $A$ [^33] A topological space possessing property (ii) above is called Fréchet-Urysohn.
@@ 6.2.2. A topological result of Grothendieck
Theorem 6.2 aggregates several results in Grothendieck’s “Critères de compacité” [Gro52, Théorèmes 1 & 2, Remarque 2, Corollaire 2]. Presently, we merely offer some remarks on translating between French terms and decades-old nomenclature to their contemporary equivalents in English. Spaces $\mathrm{C}_{\mathrm{s}}(X;Y)$ (where “s” refers to the “simple” topology, i.e., of pointwise convergence) are now denoted $\mathrm{C}_{\mathrm{p}}(X;Y)$ (or just $\mathrm{C}_{\mathrm{p}}(X)$, when $Y=\mathbb{R}$). “(Relativement) semi-compact” (resp., “relativement compact”) refers to (relatively) countably compact (resp., relatively compact) sets. Functions take values in $Y$, which we take to be a Tychonoff space (“complètement régulier”—i.e., completely regular and Hausdorff in the standard contemporary sense) endowed with an embedding into a product $\mathbb{R}^{\mathcal{P}}$, hence $Y$ is a uniform Hausdorff space (“espace uniforme séparé”[Eng89, Sections §1.5, §3.10, §8.1]. ∎
@@ 6.2.3. The Fundamental Theorem of Definability
    One could take a probabilistic approach to the uniqueness and computability of equilibria inspired by ideas from deep learning and the Examples 5.5 and 5.6 in Section 5.3. For simplicity, assume that $L$ is realcompact (so $L=\mathfrak{L}_{\mathrm{Sh}}=\mathfrak{L}$). The uniqueness and continuity of deep iterates $\gamma^{*}$ at a state $v\in L$ may be tested empirically by taking finitely many independent random points $(v_{i})_{i<k}$ in a small neighborhood of $v$ and computing $w_{i}=\gamma^{n_{i}}(v_{i})$ for some large and also random integers $(n_{i})_{i<k}$. To the extent that the points $(w_{i})_{i<k}$ are (or are not) near each other, one may infer (in a statistical sense) whether $f^{*}$ is (or is not) continuous at $v$ with increasingly larger probability as $k$ grows. At points of continuity $v$ (as determined with high probability taking $k$ sufficiently large), any of the computed points $w_{i}$ may be regarded as an approximation to the exact and unique value $\gamma^{*}(v)$. This approach hints at a relativized notion of computability based on almost-everywhere (or at least local) continuity rather than everywhere continuity, which we intend to revisit in a sequel paper.
@@ A.1.1. Unique Deep Equilibria
An empirical observation in the context of Neural Network Deep Equilibrium Models [BKK19] is that, in situations where a deep iterate $\gamma^{\mathcal{U}}=\operatorname{\mathcal{U}lim}_{n}\gamma^{n}$ of some computation $\gamma$ (assumed confined, for simplicity) exists, it is often independent of the ultrafilter $\mathcal{U}$ [^36] Implicitly, both [CRBD18] and [BKK19] work in a setting where the states space $L=\mathfrak{L}$ is realcompact, so there is no distinction between transforms and transitions-in-type. In such case, all deep iterates $\gamma^{\mathcal{U}}$ are one and the same transition $\gamma^{*}:L\to L$—a “deep state” of the NN obtained by iteration of $\gamma$. Therefore, the sequence of iterates $\gamma^{n}$ converges pointwise to the t-t $\gamma^{*}$ as an ordinary limit (rather than only as an ultralimit). We say that such $\gamma$ has the Unique Deep Equilibrium (UDE) Property. Smoothness properties of $\gamma$ are required for important applications, as described below.
@@ A.1.2. Fixed-point algorithms as “Black Boxes”
Bai et al. note (empirically) that NNs obtained by iterating a common “weight-tied” layer transition $\gamma$, the deep state $\gamma^{*}$ takes any input state $v\in L$ to another $v^{*}=\gamma^{*}(v)$ that is fixed by (the t-t implied by) the original $\gamma$, i.e., $\gamma(v^{*})=v^{*}$; in other words, $\gamma^{*}$ takes values in the set Fix$(\gamma)=\{v\in L:\gamma(v)=v\}$, so $\gamma^{*}$ is a deep equilibrium (DEQ) in a very strong sense. Empirical findings also suggest that, given $\gamma\in\Gamma$, the DEQ state $\gamma^{*}:L\to\mathrm{Fix}(\gamma)\subseteq L$ may be well approximated by some generic “black-box” fixed point algorithm $\mathtt{FindFix}$. Such an algorithm should take as inputs the transformation $\gamma$ and initial state-in-type $v$, and returns the fixed point $v^{*}=\gamma^{*}(v)=\mathtt{FindFix}(\gamma;v)$.
@@ A.1.4. Training deep networks
Training the deep neural network $F^{*}_{x}$ translates to finding weights $x$ such that $F^{*}_{x}$ satisfies a given condition, which we presently take to mean minimizing a given/specified real-valued loss function $\ell:\mathfrak{T}\to[0,\infty)$. (At least intuitively, if not necessarily literally, the value $\ell(\mathfrak{g})\geq 0$ captures how far a transition $\mathfrak{g}\in\mathfrak{T}$ is from an optimal/idealized $G\in\mathfrak{T}$.) Regarding $F^{*}_{x}$ for fixed $x$ as implicitly defined by either a fixed-point condition or ODE as above, the enormous memory cost of back-propagation through layers [^37] Not least, because back-propagation would involve an unbounded number of ordinary layers to begin! is replaced by that of minimizing the function ~$\tilde{\ell}\coloneqq\ell\circ F^{*}:X\to[0,\infty)$. Note that $\tilde{\ell}$ is merely a new real-valued predicate on the parameters CSS $\underline{X}$. Assuming that $\tilde{\ell}$ is shard-continuous, it is definable, hence depends de facto on only finitely many features $Q_{1},\dots,Q_{n}\in\mathcal{Q}$ of its input $x\in X$ (up to an arbitrarily small admissible error $\varepsilon>0$). Assuming $\tilde{\ell}$ is smooth as well, the deep network may be trained using standard/“black-box” gradient-based procedures to find a minimizer $x\in X$ for $\tilde{\ell}$. However, we note that it is essential for $\tilde{\ell}$ to be definable in order to allow even the possibility that some algorithm involving floating-point arithmetic and finitely many real quantities at a time succeeds in finding the minimizer.
@@ A.2. Neural ODEs à la Chen-Rubanova-Bettencourt-Duvenaud
where $\mathbf{s}:L\times[0,+\infty)\to\mathrm{T}L\subseteq\mathbb{R}^{\mathcal{P}}$ is a section of the tangent space $\mathrm{T}L$ of the layer state space $L$ (i.e., $\mathbf{s}(v;t)\in\mathrm{T}_{v}L$ for all $v\in L$ and $t\geq 0$, where $\mathrm{T}_{v}L$ is the tangent space of $L$ at $v$).Interpreting the ODE (A.1) hinges on the smooth manifold structure assumed of the state space $L$ [^38] The notions of differentiable structure and tangent space on an arbitrary layer space $L$ are neither well nor uniquely defined when $L$ is not finite-dimensional; their formalization would require much stronger assumptions on $L$, as well as the formalism of Banach spaces for tangent spaces $\mathrm{T}_{v}L$. Chen et al. illustrate empirically the feasibility and effectiveness of modeling deep equilibria by Neural ODEs. Let us denote the time-$t$ evolution by (A.1) using the (hopefully, suggestive) notation $v\mapsto e^{t\mathbf{s}}(v)$, i.e., $e^{t\mathbf{s}}$ is the deep equilibrium of the Neural ODE $e^{t\mathbf{s}})$ solving (A.1) (i.e., “$\gamma^{(t)}$” in the earlier informal discussion). [^39] When the section $\mathbf{s}=\mathbf{s}(v)$ depends only on the state $v$ (not on time $t$), the ODE (A.1) is autonomous. A time evolution $e^{t\mathbf{s}}$ of such autonomous Neural ODE is analogous to a “parameter-tied” deep equilibrium after Bai et al. Effective computation of $e^{t\mathbf{s}}$ relies on a generic “black-box” ODE solver algorithm $\mathtt{ODEsolve}$. Such algorithm should take as inputs the section $\mathbf{s}$, initial state-in-type $v$ and time $t\geq 0$, and returns the output $e^{t\mathbf{s}}(v)=\mathtt{ODEsolve}(\mathbf{s};v,t)$. (More realistically, such $\mathtt{ODEsolve}$ presumably would return approximate values for any finitely many specified features of $e^{t\mathbf{s}}(v)$; refer to the discussion of $\mathtt{FindFix}_{k}$ above.)

### v5

@@ 5.2.1. Relative compacta of continuous layer transitions
For any topological space $X$, let $\mathrm{C}_{\mathrm{p}}(X)\subseteq\mathbb{R}^{X}$ be the set of all continuous real functions on $X$, endowed with the relative (subspace) topology of the product $\mathbb{R}^{X}$, i.e., the topology of point-wise convergence at each $x\in X$. More generally, given two spaces $X,Y$, the space $\mathrm{C}_{\mathrm{p}}(X;Y)$ is the subspace of the product $Y^{X}=\prod_{x\in X}Y$ consisting of continuous maps $X\to Y$. ($\mathrm{C}_{\mathrm{p}}$” means “pointwise topology on continuous functions”.)
@@
A topological space $Y$ is angelic if (i) every relatively countably compact subset $A\subseteq Y$ is relatively compact, and (ii) the closure $\overline{A}\subseteq Y$ of any such (relatively compact) $A$ consists precisely of limits of sequences in AA. [^31] A topological space possessing property (ii) above is called Fréchet-Urysohn.
@@ 5.2.2. A topological result of Grothendieck
Theorem 5.2 aggregates several results in Grothendieck’s “Critères de compacité” [Gro52, Théorèmes 1 & 2, Remarque 2, Corollaire 2]. Presently, we merely offer some remarks on translating between French terms and decades-old nomenclature to their contemporary equivalents in English. Spaces $\mathrm{C}_{\mathrm{s}}(X;Y)$ (where “s” refers to the “simple” topology, i.e., of pointwise convergence) are now denoted $\mathrm{C}_{\mathrm{p}}(X;Y)$ (or just $\mathrm{C}_{\mathrm{p}}(X)$, when $Y=\mathbb{R}$). “(Relativement) semi-compact” (resp., “relativement compact”) refers to (relatively) countably compact (resp., relatively compact) sets. Functions take values in $Y$, which we take to be a Tychonoff space (“complètement régulier”—i.e., completely regular and Hausdorff in the standard contemporary sense) endowed with an embedding into a product $\mathbb{R}^{\mathcal{P}}$, hence $Y$ is a uniform Hausdorff space (“espace uniforme séparé”) [Eng89, Sections §1.5, §3.10, §8.1]. ∎
@@ 5.2.3. The Fundamental Theorem of Definability
    One could take a probabilistic approach to the uniqueness and computability of equilibria inspired by ideas from deep learning and the Examples 4.5 and 4.6 in Section 4.3. For simplicity, assume that $L$ is realcompact (so $L=\mathfrak{L}_{\mathrm{Sh}}=\mathfrak{L}$). The uniqueness and continuity of deep iterates $\gamma^{*}$ at a state $v\in L$ may be tested empirically by taking finitely many independent random points $(v_{i})_{i<k}$ in a small neighborhood of $v$ and computing $w_{i}=\gamma^{n_{i}}(v_{i})$ for some large and also random integers $(n_{i})_{i<k}$. To the extent that the points $(w_{i})_{i<k}$ are (or are not) near each other, one may infer (in a statistical sense) whether $f^{*}$ is (or is not) continuous at $v$ with increasingly larger probability as $k$ grows. At points of continuity $v$ (as determined with high probability taking $k$ sufficiently large), any of the computed points $w_{i}$ may be regarded as an approximation to the exact and unique value $\gamma^{*}(v)$. This approach hints at a relativized notion of computability based on almost-everywhere (or at least local) continuity rather than everywhere continuity, which we intend to revisit in a sequel paper.
@@ A.1.1. Unique Deep Equilibria
An empirical observation in the context of Neural Network Deep Equilibrium Models [BKK19] is that, in situations where a deep iterate $\gamma^{\mathcal{U}}=\operatorname{\mathcal{U}lim}_{n}\gamma^{n}$ of some computation $\gamma$ (assumed confined, for simplicity) exists, it is often independent of the ultrafilter $\mathcal{U}$. [^34] Implicitly, both [CRBD18] and [BKK19] work in a setting where the states space $L=\mathfrak{L}$ is realcompact, so there is no distinction between transforms and transitions-in-type. In such case, all deep iterates $\gamma^{\mathcal{U}}$ are one and the same transition $\gamma^{*}:L\to L$—a “deep state” of the NN obtained by iteration of $\gamma$. Therefore, the sequence of iterates $\gamma^{n}$ converges pointwise to the t-t $\gamma^{*}$ as an ordinary limit (rather than only as an ultralimit). We say that such $\gamma$ has the Unique Deep Equilibrium (UDE) Property. Smoothness properties of $\gamma$ are required for important applications, as described below.
@@ A.1.2. Fixed-point algorithms as “Black Boxes”
[…]
@@ A.1.4. Training deep networks
Training the deep neural network $F^{*}_{x}$ translates to finding weights $x$ such that $F^{*}_{x}$ satisfies a given condition, which we presently take to mean minimizing a given/specified real-valued loss function $\ell:\mathfrak{T}\to[0,\infty)$. (At least intuitively, if not necessarily literally, the value $\ell(\mathfrak{g})\geq 0$ captures how far a transition $\mathfrak{g}\in\mathfrak{T}$ is from an optimal/idealized $G\in\mathfrak{T}$.) Regarding $F^{*}_{x}$ for fixed $x$ as implicitly defined by either a fixed-point condition or ODE as above, the enormous memory cost of back-propagation through layers [^35] Not least, because back-propagation would involve an unbounded number of ordinary layers to begin! is replaced by that of minimizing the function $\tilde{\ell}\coloneqq\ell\circ F^{*}:X\to[0,\infty)$. Note that $\tilde{\ell}$ is merely a new real-valued predicate on the parameters CSS $\underline{X}$. Assuming that $\tilde{\ell}$ is shard-continuous, it is definable, hence depends de facto on only finitely many features $Q_{1},\dots,Q_{n}\in\mathcal{Q}$ of its input $x\in X$ (up to an arbitrarily small admissible error $\varepsilon>0$). Assuming $\tilde{\ell}$ is smooth as well, the deep network may be trained using standard/“black-box” gradient-based procedures to find a minimizer $x\in X$ for $\tilde{\ell}$. However, we note that it is essential for $\tilde{\ell}$ to be definable in order to allow even the possibility that some algorithm involving floating-point arithmetic and finitely many real quantities at a time succeeds in finding the minimizer.
@@ A.2. Neural ODEs à la Chen-Rubanova-Bettencourt-Duvenaud
where $\mathbf{s}:L\times[0,+\infty)\to\mathrm{T}L\subseteq\mathbb{R}^{\mathcal{P}}$ is a section of the tangent space $\mathrm{T}L$ of the layer state space $L$ (i.e., $\mathbf{s}(v;t)\in\mathrm{T}_{v}L$ for all $v\in L$ and $t\geq 0$, where $\mathrm{T}_{v}L$ is the tangent space of $L$ at $v$).Interpreting the ODE (A.1) hinges on the smooth manifold structure assumed of the state space LL. [^36] The notions of differentiable structure and tangent space on an arbitrary layer space $L$ are neither well nor uniquely defined when $L$ is not finite-dimensional; their formalization would require much stronger assumptions on $L$, as well as the formalism of Banach spaces for tangent spaces $\mathrm{T}_{v}L$. Chen et al. illustrate empirically the feasibility and effectiveness of modeling deep equilibria by Neural ODEs. Let us denote the time-$t$ evolution by (A.1) using the (hopefully, suggestive) notation $v\mapsto e^{t\mathbf{s}}(v)$, i.e., $e^{t\mathbf{s}}$ is the deep equilibrium of the Neural ODE $e^{t\mathbf{s}})$ solving (A.1) (i.e., $\gamma^{(t)}$” in the earlier informal discussion). [^37] When the section $\mathbf{s}=\mathbf{s}(v)$ depends only on the state $v$ (not on time $t$), the ODE (A.1) is autonomous. A time evolution $e^{t\mathbf{s}}$ of such autonomous Neural ODE is analogous to a “parameter-tied” deep equilibrium after Bai et al. Effective computation of $e^{t\mathbf{s}}$ relies on a generic “black-box” ODE solver algorithm $\mathtt{ODEsolve}$. Such algorithm should take as inputs the section $\mathbf{s}$, initial state-in-type $v$ and time $t\geq 0$, and returns the output $e^{t\mathbf{s}}(v)=\mathtt{ODEsolve}(\mathbf{s};v,t)$. (More realistically, such $\mathtt{ODEsolve}$ presumably would return approximate values for any finitely many specified features of $e^{t\mathbf{s}}(v)$; refer to the discussion of $\mathtt{FindFix}_{k}$ above.)